What sample size is needed to give a margin of error within plus-or-minus 2.5% in estimating a population proportion with 90% confidence? An initial small sample has p ^ =equals 0.81. Round your answer up to the nearest integer. The absolute tolerance is +/-2
Answer #1

Solution :

Given that,

\hat p = 0.81

1 - \hat p = 1 - 0.81 = 0.19

margin of error = E = 0.025

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p )

= (1.645 /0.025)2 * 0.81* 0.19

=666

Sample size = 666